OpenStudy (bananas1234):

Find the distance between point A (-3, 5) and point B (5, -7) in the coordinate plane.

pooja195 (pooja195):

Use the distance formula \[\huge d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\]

OpenStudy (bananas1234):

ok so like (5^2 - (-3)^1) + (-7^2 - 5^1)^2?

pooja195 (pooja195):

\[\huge d=\sqrt{(5--3)^2+(-7-5)^2}\] \[\huge d=\sqrt{(5+3)^2+(-7-5)^2}\] 5+3=?

OpenStudy (bananas1234):

8

pooja195 (pooja195):

\[\huge d=\sqrt{(8)^2+(-7-5)^2}\] -7-5=?

OpenStudy (bananas1234):

-12

pooja195 (pooja195):

\[\huge d=\sqrt{(8)^2+(-12)^2}\] 8^2=?

OpenStudy (bananas1234):

64

pooja195 (pooja195):

\[\huge d=\sqrt{64+(-12)^2}\] -12^2=?

OpenStudy (bananas1234):

-144

pooja195 (pooja195):

it would be just 144 not -144 because -12 x -12=144 (neg *neg=pos) \[\huge d=\sqrt{64+144}\] 144+64=?

OpenStudy (bananas1234):

208

pooja195 (pooja195):

Good so \[\huge~\rm~\bf~\sqrt{208}~or~14.42\]

pooja195 (pooja195):

Depending on what your question asks for.

OpenStudy (bananas1234):

awesome thanks, that helped alot :)

pooja195 (pooja195):

yw ^_^